Driving 1000 miles a month is not unusual for a short-distance commuter. If your vehicle gets 26.70 mpg, you would use approximately 37.45 gallons of gasoline every month. If gasoline is approximated as C8H18 (density = 0.703 g/mL), how many grams of carbon dioxide does your vehicle emit every month?
1
Expert's answer
2018-03-27T08:02:13-0400
Solution:
37.45 gallons is 141.7637 L = 141763.7 mL
ρ = m / V, so m (octane) = ρV = 0.703 g/mL x 141763.7 mL = 99659.8811 g
2C8H18 + 25O2 = 16CO2 + 18H2O
M (C8H18) = 114.22852 g/mol
M (CO2) = 44.009 g/mol
2 moles (2x114.22852 g) of octane produce 16 moles (16x44.009 g) of CO2 gas
99659.8811 g produce x g
x = 99659.8811 g x 16x44.009 g / 2x114.22852 g = 307168.94 g
Answer: 307168.94 g.